# MUSE: Mamba is Efficient Multi-scale Learner for Text-video Retrieval

### v1

@@ Abstract
Text-Video Retrieval (TVR) aims to align and associate relevant video content with corresponding natural language queries. Most existing TVR methods are based on large-scale pre-trained vision-language models (e.g., CLIP). However, due to CLIP’s inherent plain structure, few TVR methods explore the multi-scale representations which offer richer contextual information for a more thorough understanding. To this end, we propose MUSE, a multi-scale mamba with linear computational complexity for efficient cross-resolution modeling. Specifically, the multi-scale representations are generated by applying a feature pyramid on the last single-scale feature map. Then, we employ the Mamba structure as an efficient multi-scale learner to jointly learn scale-wise representations. Furthermore, we conduct comprehensive studies to investigate different model structures and designs. Extensive results on three popular benchmarks have validated the superiority of MUSE.  [^1] The code is available at https://github.com/hrtang22/MUSE.
@@
Introduction
Text-Video Retrieval (TVR) (Gabeur et al. 2020; Gorti et al. 2022; He et al. 2021; Lei et al. 2021; Luo et al. 2022; Ma et al. 2022; Wang et al. 2022; Yang, Cao, and Zou 2023; Cao et al. 2024) is a fundamental task in multimodal research. Its objective is to locate the most relevant video content within a repository in response to a text query and vice versa.
Based on large-scale image-text pre-trained model CLIP (Radford et al. 2021), most current TVR methods focus on transferring CLIP to the video-text domain. To achieve fine-grained representations, mainstream methods capture cross-modal alignment at different granularities, including video-sentence (Ma et al. 2022), frame-sentence (Gorti et al. 2022; Ma et al. 2022) or even patch-word (Wang et al. 2023) levels. However, CLIP is inherently a plain structure with the identical token length for all the layers. Therefore, these methods ignore the exploration of representations of different scales [^2] We use “scale” for feature-level representation and “resolution” for pixel-level raw image., which provides more valuable contextual information for comprehensive understanding. For example, in Figure 1(a), the textual query aims to retrieve a video where “people are carrying torches and chasing a giant squidward”. As shown, the most discriminative object torch is not highlighted in the original resolution. The loss of such detailed yet important information leads to incorrect retrieval results. In contrast, we can see that the torch region is correctly highlighted when focusing on higher-resolution representations. Therefore, information hidden in higher resolution should be considered for TVR.
@@ Introduction
Text-Video Retrieval. TVR (Yu, Kim, and Kim 2018; Gabeur et al. 2020; Gorti et al. 2022; He et al. 2021; Lei et al. 2021; Luo et al. 2022; Ma et al. 2022; Wang et al. 2022; Cao et al. 2024; Jiang et al. 2022; Jin et al. 2023b, 2022, c; Mao et al. 2023; Wu et al. 2024) is a pivotal task for video cross-modal learning, which has widespread application in video understanding (Liu et al. 2024c; Zhang et al. 2021, 2022; Li et al. 2023a; Cao et al. 2022c, a; Li et al. 2023b; Cao et al. 2023) and multi-modal interactions (Cao et al. 2021, 2022b; Liu et al. 2023a; Ye et al. 2023; Ji et al. 2023; Luo et al. 2024; Liu et al. 2024b). With the advancement of image-text pretraining, recent works (Luo et al. 2022; Ma et al. 2022; Gorti et al. 2022; Jiang et al. 2022; Jin et al. 2023b) resort to image-text pretraining model CLIP (Radford et al. 2021) and focus on image-to-video transferring fine-tuning. The primary work CLIP4clip (Luo et al. 2022) investigates three kinds of temporal aggregation manners, which enlightens the follow-up works. To achieve fine-grained representations, X-CLIP (Ma et al. 2022) explores cross-grained contrastive learning, including video-sentence, video-word, frame-word and frame-sentence. Hunyuan_tvr (Jiang et al. 2022) divides the video-language interaction into frame-word, clip-phrase, and video-sentence granularities. DiCoSA (Jin et al. 2023b) further improves the fine-grained alignments by disentangling video into visual concepts. However, these methods neglect the cross-resolution relationships, which offer another perspective over the resolution-wise feature correlations. Our proposed MUSE bridges this gap by presenting scale-aware representations.
@@
Mamba for Video. Based on the success of Mamba in language modeling, Vim (Zhu et al. 2024) and VMamba (Liu et al. 2024d) has pioneered Mamba architecture in vision by designing a bidirectional State Space Model and 2D selective scan manners for image recognition. The follow-up methods (Yang et al. 2024; Huang et al. 2024; Pei, Huang, and Xu 2024) explore different architecture design and selective scan manners. Also, some works (Shen et al. 2024; Ma, Li, and Wang 2024; Liu et al. 2024a; Chen et al. 2024b) have extended the success of Mamba to areas such as medical image processing, 3D reconstruction, point cloud understanding, etc. For video understanding, VideoMamba (Li et al. 2024) first trains a video foundation model with Mamba backbone using the paradigm of unmasked teacher (Li et al. 2023c) and shows its efficiency compared with the transformer counterpart. Video mamba suite (Chen et al. 2024a) explores Mamba’s effectiveness in the video downstream tasks by replacing attention blocks with its DBM blocks. In this work, we built a Mamba learner with a simple gated structure based on the Bidirectional Mamba (Zhu et al. 2024). We find that Mamba is an efficient multi-scale video learner that surpasses Transformer and other linear attention methods in text-video retrieval.
@@
$\boldsymbol{v}^{a}_{l+1}$ $=\mathcal{G}(\operatorname{Norm}(\boldsymbol{y}_{l}))+\boldsymbol{y}_{l},$
@@
Table 1: Plut-and-play experiments on MSR-VTT. We compare the text-to-video retrieval results before and after adding our proposed method, MUSE, on four baseline models. $\dagger$ denotes our reproduction of the method.
@@
Why using Mamba? To answer the question of why Mamba is our model selection, we testify the superiority of Mamba in two perspectives: (1) Memory Usage. We compare the memory usage between Transformer attention of quadratic complexity and Mamba of Linear complexity by calculating the memory usage as the input frame number grows. As shown in Figure 3, we test the memory usage based on CLIP4clip with batch size 16 and scale {1, 3, 7, 14}. The green line reflects the memory usage of the base model using mean pooling without applying MUSE. The pink line with marker star and blue line with dot reflects the memory growth when leveraging Mamba and Transformer as the video learner after multi-scale feature extraction and aggregation. From the comparison, we can easily identify that Mamba has a remarkably lower computation resource requirement than Transformer attention. For instance, when the input frame number comes to 20, the model’s memory with Transformer needs more than 80GB GPU memory on a single GPU, which exceeds the memory-bound of most modern hardware accelerators. In contrast, Mamba only needs 15.37GB memory at input frame 16, which saves 79.7% (76.34GB) memory resources of the Transformer. Due to the linear complexity of Mamba, the memory growth is more acceptable even when it comes to 64 frames input (47.20GB in memory). (2) Performance. To further validate the effectiveness of Mamba architecture, we compare the performance of Mamba with Transformer and other models with linear computational complexity including MambaOut (Yu and Wang 2024) and FlashAttention (Dao et al. 2022). As shown in Table 4, Mamba performs best on MST-VTT among the four listed models. We replace Mamba block with other counterparts for a fair comparison. Compared with FlashAttention and Transformer, Mamba achieves better performance with almost the same memory usage as FlashAttention. To validate the effectiveness of SSM architecture in Mamba, we compare it with MambaOut which is implemented by only removing the SSM module. Results show that most of the ability to model video sequences comes from the SSM structure of Mamba.
@@
Figure 3: Comparison of the memory usage among Transformer, Mamba, Baseline. The baseline selected is CLIP4clip(Luo et al. 2022) with mean pooling for feature aggregation.
Table 5: Ablations on scan strategies.
@@
Table 6: Ablations on aggregation manners.

### v2

@@ Abstract
Text-Video Retrieval (TVR) aims to align and associate relevant video content with corresponding natural language queries. Most existing TVR methods are based on large-scale pre-trained vision-language models (e.g., CLIP). However, due to CLIP’s inherent plain structure, few TVR methods explore the multi-scale representations which offer richer contextual information for a more thorough understanding. To this end, we propose MUSE, a multi-scale mamba with linear computational complexity for efficient cross-resolution modeling. Specifically, the multi-scale representations are generated by applying a feature pyramid on the last single-scale feature map. Then, we employ the Mamba structure as an efficient multi-scale learner to jointly learn scale-wise representations. Furthermore, we conduct comprehensive studies to investigate different model structures and designs. Extensive results on three popular benchmarks have validated the superiority of MUSE.
Codehttps://github.com/hrtang22/MUSE
Extended versionhttps://arxiv.org/abs/2408.10575
Introduction
Text-Video Retrieval (TVR) (Gabeur et al. 2020; Gorti et al. 2022; He et al. 2021; Lei et al. 2021; Luo et al. 2022; Ma et al. 2022; Wang et al. 2022) is a fundamental task in multimodal research. Its objective is to locate the most relevant video content within a repository in response to a text query and vice versa.
Based on large-scale image-text pre-trained model CLIP (Radford et al. 2021), most current TVR methods focus on transferring CLIP to the video-text domain. To achieve fine-grained representations, mainstream methods capture cross-modal alignment at different granularities, including video-sentence (Ma et al. 2022), frame-sentence (Gorti et al. 2022; Ma et al. 2022) or even patch-word (Wang et al. 2023) levels.
However, CLIP is inherently a plain structure with the identical token length for all the layers. Therefore, these methods ignore the exploration of representations of different scales, which provides more valuable contextual information for comprehensive understanding. For example, in Figure 1(a), the textual query aims to retrieve a video where “people are carrying torches and chasing a giant squidward”. As shown, the most discriminative object torch is not highlighted in the original resolution [^1] We use “scale” for feature-level representation and “resolution” for pixel-level raw image.. The loss of such detailed yet important information leads to incorrect retrieval results. In contrast, we can see that the torch region is correctly highlighted when focusing on higher-resolution representations. Therefore, information hidden in higher resolution should be considered for TVR.
@@ Introduction
Text-Video Retrieval. TVR (Yu, Kim, and Kim 2018; Gabeur et al. 2020; Lei et al. 2020; Gorti et al. 2022; He et al. 2021; Lei et al. 2021; Luo et al. 2022; Ma et al. 2022; Wang et al. 2022; Cao et al. 2024; Jiang et al. 2022; Jin et al. 2023b, 2022, c; Wu et al. 2024) is a pivotal task for video cross-modal learning, which has widespread application in video understanding (Liu et al. 2024c; Zhang et al. 2021, 2022; Li et al. 2023a; Cao et al. 2022c, a; Li et al. 2023b; Cao et al. 2023, 2022a) and multi-modal interactions (Cao et al. 2021, 2022b; Liu et al. 2023a; Ye et al. 2023; Ji et al. 2023; Luo et al. 2024; Liu et al. 2024b; Yang et al. 2021; Yang, Cao, and Zou 2023; Li et al. 2024a). With the advancement of image-text pretraining, recent works (Luo et al. 2022; Ma et al. 2022; Tu et al. 2022; Gorti et al. 2022; Jiang et al. 2022; Jin et al. 2023b) resort to image-text pretraining model CLIP (Radford et al. 2021) and focus on image-to-video transferring fine-tuning. The primary work CLIP4clip (Luo et al. 2022) investigates three kinds of temporal aggregation manners, which enlightens the follow-up works. To achieve fine-grained representations, X-CLIP (Ma et al. 2022) explores cross-grained contrastive learning, including video-sentence, video-word, frame-word and frame-sentence. Hunyuan_tvr (Jiang et al. 2022) divides the video-language interaction into frame-word, clip-phrase, and video-sentence granularities. DiCoSA (Jin et al. 2023b) further improves the fine-grained alignments by disentangling video into visual concepts. However, these methods neglect the cross-resolution relationships, which offer another perspective over the resolution-wise feature correlations. Our proposed MUSE bridges this gap by presenting scale-aware representations.
@@
Mamba for Video. Based on the success of Mamba in language modeling, Vim (Zhu et al. 2024) and VMamba (Liu et al. 2024d) has pioneered Mamba architecture in vision by designing a bidirectional State Space Model and 2D selective scan manners for image recognition. The follow-up methods (Yang et al. 2024; Huang et al. 2024; Pei, Huang, and Xu 2024) explore different architecture design and selective scan manners. Also, some works (Shen et al. 2024; Ma, Li, and Wang 2024; Liu et al. 2024a; Chen et al. 2024b) have extended the success of Mamba to areas such as medical image processing, 3D reconstruction, point cloud understanding, etc. For video understanding, VideoMamba (Li et al. 2024b) first trains a video foundation model with Mamba backbone using the paradigm of unmasked teacher (Li et al. 2023c) and shows its efficiency compared with the transformer counterpart. Video mamba suite (Chen et al. 2024a) explores Mamba’s effectiveness in the video downstream tasks by replacing attention blocks with its DBM blocks. In this work, we built a Mamba learner with a simple gated structure based on the Bidirectional Mamba (Zhu et al. 2024). We find that Mamba is an efficient multi-scale video learner that surpasses Transformer and other linear attention methods in text-video retrieval.
@@
$\boldsymbol{v}^{a}_{l+1}$ $=\mathcal{G}(\operatorname{Norm}(\boldsymbol{y}_{l}))+\boldsymbol{v}^{a}_{l},$
@@
Table 1: Plug-and-play experiments on MSR-VTT. We compare the text-to-video retrieval results before and after adding our proposed method, MUSE, on four baseline models. $\dagger$ denotes our reproduction of the method.
@@
Why using Mamba? To answer the question of why Mamba is our model selection, we testify the superiority of Mamba in two perspectives: (1) Memory Usage. We compare the memory usage between Transformer attention of quadratic complexity and Mamba of Linear complexity by calculating the memory usage as the input frame number grows. As shown in Figure 3, we test the memory usage based on CLIP4clip with batch size 16 and scale {1, 3, 7, 14}. The green line reflects the memory usage of the base model using mean pooling without applying MUSE. The pink line with marker star and blue line with dot reflects the memory growth when leveraging Mamba and Transformer as the video learner after multi-scale feature extraction and aggregation. From the comparison, we can easily identify that Mamba has a remarkably lower computation resource requirement than Transformer attention. For instance, when the input frame number comes to 20, the model’s memory with Transformer needs more than 80GB GPU memory on a single GPU, which exceeds the memory-bound of most modern hardware accelerators. In contrast, Mamba only needs 15.37GB memory at input frame 16, which saves 79.7% (76.34GB) memory resources of the Transformer. Due to the linear complexity of Mamba, the memory growth is more acceptable even when it comes to 64 frames input (47.20GB in memory). (2) Performance. To further validate the effectiveness of Mamba architecture, we compare the performance of Mamba with Transformer and other models with linear computational complexity including MambaOut (Yu and Wang 2024) and FlashAttention (Dao et al. 2022). As shown in Table 4, Mamba performs best on MSR-VTT among the four listed models. We replace Mamba block with other counterparts for a fair comparison. Compared with FlashAttention and Transformer, Mamba achieves better performance with almost the same memory usage as FlashAttention. To validate the effectiveness of SSM architecture in Mamba, we compare it with MambaOut which is implemented by only removing the SSM module. Results show that most of the ability to model video sequences comes from the SSM structure of Mamba.
@@
Table 5: Ablations on scan strategies.
Figure 3: Comparison of the memory usage among Transformer, Mamba, and Baseline. The baseline selected is CLIP4clip(Luo et al. 2022) with mean pooling for feature aggregation.
@@ Conclusion
Acknowledgments
This work is supported in part by National Key Research and Development Program of China (2024YFE0203100) and Guangdong Provincial Key Laboratory of Ultra High Definition Immersive Media Technology (Grant No. 2024B1212010006).
Appendix
In the appendix, we list some of the detailed illustrations of our proposed method following the suggestions of reviewers. These include illustration of multi-scale feature extraction, training time, GPU memory consumption, and illustration of ResMamba architecture.
A. Illustration of multi-scale feature extraction
In our implementation, we utilize several Conv2d and MaxPooling modules to extract multi-scale video features based on the output from CLIP, as discussed in the ”Multi-scale Feature Generation” section. For instance, if the CLIP output has dimensions 14x14 and the scale factor is s=3, we first apply a max-pooling operation to reduce the feature map to 3x3, and then use multiple Conv2d layers with LayerNorm to extract features of this scale. Through experimentation, we have found that this straightforward approach provides the most effective results, which aligns with similar strategies discussed in the ViTDet(Li et al. 2022) model.
B. Training time after adding MUSE
Regarding the time complexity, as Table 9 shows, after incorporating MUSE, the training time on the MSR-VTT dataset increased to 17.3 GPU hours, which is a 10.9% increase compared to the original CLIP4clip baseline of 15.6 GPU hours. This training time was measured using 12 video frames as input. Considering the significant performance improvement and the relatively low memory overhead, we believe that the slight increase in training time is acceptable for the community, especially for training text-video retrieval models.
Method R@1↑ R@5↑ R@10↑ Time(hr)↓
Baseline 42.6 70.8 79.9 15.6
+ MUSE 44.8 71.6 82.1 17.3
Table 9: Training time comparison after adding MUSE. The baseline model is CLIP4clip with 12 frame inputs.
C. GPU memory consumption.
We list the GPU memory consumption during training of the baseline model together with the models with MUSE or Transformer. As shown in Table 10, the memory required by MUSE is 12.6GB for each GPU when training with batch size 128 frame 12 on 8 A100s.
Method R@1↑ R@5↑ R@10↑ Memory(GB)↓
Baseline 42.6 70.8 79.9 9.84
+ MUSE 44.8 71.6 82.1 12.6 (+2.76)
+ Transformer 43.0 71.1 80.0 36.8 (+26.96)
Table 10: GPU memory comparison. We list the memory consumption of the Baseline model (CLIP4clip), MUSE, and its Transformer counterpart.
D. Detailed illustration of ResMamba architecture
In our paper, we advance Mamaba to ResMamba with the following two considerations: 1) Stable training: Residual connections have been widely used in various tasks for stabling gradient back-propagation; 2) Feature retaining: This simple skip-connection maintains the un-traversed features in the fused features. Specifically, we implemented a gated network for output features to pass through before being added to the input features, it consists of a LayerNorm followed by a Linear layer with zero initialization. The ablations are shown in Table 11.
Method R@1↑ R@5↑ R@10↑ MnR↓
w/o residual 43.9 71.7 81.6 15.1
w/ residual 44.8 71.6 82.1 15.6
Table 11: Ablations of residual architecture.